Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below:

  

Sales (13,400 units × $20 per unit) $ 268,000
Variable expenses 160,800
Contribution margin 107,200
Fixed expenses 119,200
Net operating loss $ (12,000 )

Required:

1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales.

2. The president believes that a $6,000 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will increase unit sales and the total sales by $89,000 per month. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income?

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $32,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)?

4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by $0.70 per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,700?

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $60,000 each month.

a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales.

b. Assume that the company expects to sell 20,900 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.)

c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,900 units)?

Compute the company’s CM ratio and its break-even point in unit sales and dollar sales. (Do not round intermediate calculations. Round "CM ratio" to the nearest whole percentage (i.e., 0.234 should be entered as "23").

CM ratio %
Break-even point in unit sales
Break-even point in dollar sales

The president believes that a $6,000 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will increase unit sales and the total sales by $89,000 per month. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income? (Do not round intermediate calculations.)

by

Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $32,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)? (Losses should be entered as a negative value.)

Revised net operating income (loss)

Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by $0.70 per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,700? (Do not round intermediate calculations. Round final answer to the nearest whole unit.)

Show less

Unit sales to attain target profit

Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $60,000 each month. Compute the new CM ratio and the new break-even point in unit sales and dollar sales. (Do not round intermediate calculations. Round "CM ratio" to the nearest whole percentage (i.e., 0.234 should be entered as "23") and other answers to the nearest whole number.)

Show less

CM ratio %
Break-even point in unit sales
Break-even point in dollar sales
  • Req
Answer #1

1. CM ratio and break even point

CM ratio = Contribution margin / sales

Contribution margin = $107,200

Sales = $268,000

CM ratio = 107,200 / 268,000 = 0.4

CM ratio in percentage = 0.4 * 100 = 40%

Break even point in unit sales = Fixed cost / contribution margin per unit

Break even point in dollar sales = Fixed cost / contribution margin ratio

Fixed cost = $119,200

Contribution margin per unit = Total contribution margin / number of units sold

Number of units sold = 13,400 units

Contribution margin per unit = 107,200 / 13,400 = $8

Break even point in unit sales = 119,200 / 8 = 14,900 units

Break even point in dollar sales = 119,200 / 0.4 = $298,000

2.Increase in operating income after advertising of $6,000 adopted

Working note Amount
Sales 268,000 + 89,000 additional sales $357,000

Less variable cost

Total variable cost in question = $160,800

Number of units sold in question = 13,400

Variable cost per unit = total variable cost / unit sales = 160,800 / 13,400 = $12

At present, unit sales = present sales / selling price

= 357,000 / 20 = 17,850 units

Total variable cost = unit sales * variable cost per unit

= 17,850 * 12 = $214,200

($214,200)
Contribution margin $142,800
Less Fixed cost $119,200 + $6,000 (for advertisement) ($125,200)
Net operating income $17,600

Before advertising, the income was -$12,000(loss) and after advertising , it is $17,600. The difference is $29,600 and that is the increase.

3. Net operating income/loss under new policy of sales manager

unit sales policy doubles the unit sales= 13,400*2 26,800 units
Selling price 10% reduction = $20*90% $18
Total sales in dollars Unit sales * selling price= 26,800 * 18 $482,400
Less variable cost Unit sales* variable cost per unit = 26,800*12 ($321,600)
Contribution margin $160,800
Less fixed cost $119,200 + $32,000 (advertising) ($151,200)
Net operating income $9,600

4. Target unit sales to achieve profit of $4,700

Target unit sales = (fixed cost + desired profit) / contribution margin per unit

Fixed cost $119,200

Contribution margin per unit

= selling price - variable cost per unit

= 20 - (12 + 0.7 for additional packing cost)= 20 - 12.7

$7.3
Desired profit $4,700

Target unit sales = (119,200 + 4,700) / 7.3 = 123,900 / 7.3 = 16,972.6 rounded to 16,973 units

5.a) New CM ratio

Selling price = $20, Variable cost = 12 - 3  (reduction) = 9

Contribution margin per unit = Sellin price - variable cost = 20 - 9 = $11

CM ratio = contribution margin per unit / selling price = 11/20 = 0.55

Break even point in unit sales = Fixed cost / contribution margin per unit

Fixed cost = 119,200 + 60,000(additional) = $179,200

Break even point in unit sales = 179,200 / 11 = 16,291 units

Break even point in dollar sales = fixed cost / contribution margin ratio = 179,200 / 0.55 =$325,818

5.b) Income statement if operations are automated

units= 20,900

PEM, INC
INCOME STATEMENT
Per unit % Total
Sales $20 100% $418,000
Less variable cost ($9) 45% ($188,100)
Contribution margin $11 55% $229,900
Less fixed cost ($179,200)
Profit $50,700

income statement if operations are not automated

units = 20,900

PEC,INC
INCOME STATEMENT
Per unit % Total
Sales $20 100% $418,000
Less variable cost ($12) 60% ($250,800)
Contribution margin $8 40% $167,200
Less fixed cost ($119,200)
Profit $48,000

5.c. Yes, the operations can be automated since it brings an increased profit. The profit has increased from $48,000 to $50,700. The incremental profit due to automation is $2,700.

(2,700 = 50,700 - 48,000)